OpenStudy (anonymous):

sin theta =-4/5 cos beta=-5/13 Find the exact value of tan(theta+beta) Will FAN and MEDAL

OpenStudy (anonymous):

easy -1 12/65

OpenStudy (anonymous):

or answer:= -77/65= -1 12/65

OpenStudy (alyssa_xo):

bit of a horse lover meself :\(\huge \color{pink} \heartsuit\)

OpenStudy (anonymous):

My aunt molested me

OpenStudy (anonymous):

lol...telu here is a song deticated to you https://www.youtube.com/watch?v=skf9JGdB4yM

OpenStudy (anonymous):

can you show me hwo you set it up? I know the formula

OpenStudy (anonymous):

@WildHorseLover2002

OpenStudy (anonymous):

how i got the whole answer

OpenStudy (anonymous):

is i added -4/5 + -5/13=77/65 then you have to Convert it to a mixed number so= 77 ÷ 65 = 1 R12 = 77 65 = 1 12 65

OpenStudy (anonymous):

so 77/65 = 1 12/65

OpenStudy (anonymous):

hope it helps

OpenStudy (anonymous):

did it help